# Supermassive black hole growth from stellar binary encounters

### v3

@@ 2 The model: black hole growth through the Hills mechanism
The Hills mechanism operates when the center of mass of a binary star comes within a distance $r_{\rm close}$ of a SMBH, where tidal forces from black hole overcome the gravitational binding of the binary pair (Hills, 1988). If $r_{\rm close}$ is within the event horizon of the SMBH (for black holes with mass $_{\bullet}\gtrsim 10^{8} M\rm M_{\odot}$), or inside the tidal radius of an individual star (for smaller black holes), the binary is certainly destroyed.
@@
Hills (1988) noted that the break-up of a binary may occur at larger $r_{\rm close}$, depending on the binary’s orbital phase and orientation. Conversely, some binaries may survive the encounter with the SMBH intact at smaller $r_{\rm close}$, again, depending on the details of their orbits. [^1] Hills (1988) framed this range of outcomes probabilistically, with the likelihood of a break-up at $r_{\rm close}=0$ being 100% and decreasing linearly to zero at a closest-approach distance of $r_{{}_{\rm D}}/a_{\rm bin}$, where $r_{{}_{\rm D}}\approx r_{\rm close}(m_{\rm bin}/11M_{\bullet})^{1/3}$. Thus, when $r_{{}_{\rm D}}$ exceeds the binary separation (a value of about twice the distance in Eq. (2)) a binary can pass by a black hole unscathed. Nonetheless, Equation (2) provides a characteristic value of the binary break-up radius (Yu & Tremaine, 2003). Below, in calculating interaction cross sections, we adopt an even smaller fiducial value of $r_{\rm close}=200$ km/s, bringing the likelihood of a break-up closer to unity.
@@
The semimajor axis of the bound partner following tidal break-up is
@@ 2.1 Binary break-up rates
where $m_{\rm avg}\approx$ 0.3 $\rm M_{\odot}$ is the average mass of a star, counting binary components as individual stars. In the Milky Way, the binary fraction is roughly half of all stars (Duchêne & Kraus, 2013, and references therein), though $f_{\rm bin}$ is dependent on the spectral type of the components (e.g., Moe & Di Stefano, 2017). Typically younger, more massive stars have a higher binary fraction than less massive, older stars.
The central mass density of stars, $\rho_{0}$, depends strongly on the SMBH mass, $M_{\bullet}$. For large elliptical galaxies, the density falls with increasing black hole mass $\rho_{0}\sim 1/M_{\bullet}$ (e.g., Faber et al., 1997). The trend is opposite in smaller galaxies. Both central stellar density and black hole mass increase with the size and mass of the galactic host. Here, we introduce the parameterization,
@@
Together, the three $nv\sigma$” terms provide an estimate of the rate of binary break-up through the Hills mechanism. In this kinematical approximation, the break-up rate of binary stars through the Hills mechanism is
@@ 3 SMBH growth in nearby galactic hosts
Figure 2 illustrates the dataset, showing the representative central stellar mass density (evaluated at 5 pc from the galactic center) versus SMBH mass for all of the sources in Hannah et al. (2024, see also ). The plot also shows our analytical expression that broadly tracks the central density as a function of black hole mass (Eq. (5)).

### v4

@@ 2 The model: black hole growth through the Hills mechanism
The Hills mechanism operates when the center of mass of a binary star comes within a distance $r_{\rm close}$ of a SMBH, where tidal forces from the black hole overcome the gravitational binding of the binary pair (Hills, 1988). If $r_{\rm close}$ is within the event horizon of the SMBH (for black holes with mass $_{\bullet}\gtrsim 10^{8} M\rm M_{\odot}$), or inside the tidal radius of an individual star (for smaller black holes), the binary is certainly destroyed.
@@
Hills (1988) noted that the break-up of a binary may occur at larger $r_{\rm close}$, depending on the binary’s orbital phase and orientation. Conversely, some binaries may survive the encounter with the SMBH intact at smaller $r_{\rm close}$, again, depending on the details of their orbits. [^1] Hills (1988) framed this range of outcomes probabilistically, with the likelihood of a break-up at $r_{\rm close}=0$ being 100% and decreasing linearly to zero at a closest-approach distance of $r_{{}_{\rm D}}/a_{\rm bin}$, where $r_{{}_{\rm D}}\approx r_{\rm close}(m_{\rm bin}/11M_{\bullet})^{1/3}$. Thus, when $r_{{}_{\rm D}}$ exceeds the binary separation (a value of about twice the distance in Eq. (2)), a binary can pass by a black hole unscathed. Nonetheless, Equation (2) provides a characteristic value of the binary break-up radius (Yu & Tremaine, 2003). Below, in calculating interaction cross sections, we adopt an even smaller fiducial value of $r_{\rm close}=200$ km/s, bringing the likelihood of a break up closer to unity.
@@
The semimajor axis of the bound partner after tidal break-up is
@@ 2.1 Binary break-up rates
where $m_{\rm avg}\approx$ 0.3 $\rm M_{\odot}$ is the average mass of a star, counting binary components as individual stars. In the Milky Way, the binary fraction is roughly half of all stars (Duchêne & Kraus, 2013, and references therein), although $f_{\rm bin}$ is dependent on the spectral type of the components (e.g., Moe & Di Stefano, 2017). Typically younger, more massive stars have a higher binary fraction than less massive, older stars.
The central mass density of stars, $\rho_{0}$, strongly depends on the SMBH mass, $M_{\bullet}$. For large elliptical galaxies, the density falls with increasing black hole mass $\rho_{0}\sim 1/M_{\bullet}$ (e.g., Faber et al., 1997). The trend is opposite in smaller galaxies. Both central stellar density and black hole mass increase with the size and mass of the galactic host. Here, we introduce the parameterization,
@@
Together, the three $nv\sigma$” factors provide an estimate of the rate of binary break-up through the Hills mechanism. In this kinematical approximation, the break-up rate of binary stars through the Hills mechanism is
@@ 3 SMBH growth in nearby galactic hosts
Figure 2 illustrates the dataset, showing the representative central stellar mass density (evaluated at 5 pc from the galactic center) versus SMBH mass for all sources in Hannah et al. (2024, see also ). The plot also shows our analytical expression that broadly tracks the central density as a function of black hole mass (Eq. (5)).